A generator at one end of a very long string creates a wave given by y = (7.44 cm) cos[(π/2)(3.80 m-1x + 3.17 s-1t)] and a generator at the other end creates the wave y = (7.44 cm) cos[(π/2)(3.80 m-1x - 3.17 s-1t)] Calculate the (a) frequency, (b) wavelength, and (c) speed of each wave. For x ≥ 0, what is the location of the node having the (d) smallest, (e) second smallest, and (f) third smallest value of x? For x ≥ 0, what is the location of the antinode having the (g) smallest, (h) second smallest, and (i) third smallest value of x?
Answer #1

a)for ist wave k=3.8

for 2nd wave k=3.8

for ist wave w=3.17

for 2nd wave w=3.17

a)frequency of each wave=w/2pi

=.5hz

b)lambda of each wave=2pi/k

=2pi/3.8

=1.65m

c)speed=w/k

=3.17/3.8

=0.83m/s